x over 30 = 3 over 15
Answer by jim_thompson5910(35256)

x/30 = 3/15

x*15 = 3*30

15x = 90

x = 90/15

x = 6


So the answer is x = 6

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim